3. Illustrate graphically

Suppose that a competitive firm's marginal cost of producing output q is given by

MC(q)= 70+6q

Assume that the market price of the firm’s product is $145.

A. At what level of output will the firm produce?

B. How much is the firm’s producer surplus?

C. Illustrate graphically profit maximization point and producer surplus.

D. Illustrate this market at a loss. Explain.
Answer #1

Given, MC = 70 + 6q
Price = $145

A) In equilibrium, P = MC
70 + 6q = 145
6q = 75
q = 75 / 6 = 12.5

Output of firm = 12.5 unit.

B) Producer surplus: 1/2 (145 - 70) *12.5 = 1/2 * 75 * 12.5 = $468.75

C) & D)

Producer surplus and profit maximising points are marked accordingly.

Since average total cost is above the equilibrium point, this signifies that at equilibrium, total revenue is lesser than total cost. Thus the market is at loss.